OpenStudy (anonymous):

What are the real and complex solutions of the polynomial equation x^3-216=0

OpenStudy (dayakar):

x^3 = 216 x^3 = 6*6*6 x^3 = 6^3 what is x value?

OpenStudy (dumbcow):

Factor this using difference of cubes \[a^3 - b^3 = (a-b)(a^2+ab+b^2)\] set each factor equal to 0, and solve using quadratic formula